Can you please explain how to get a answer to this question. Bob and Betty traveled 392 miles in 7 hours. If they travel at the same rate, how long will it take them to travel 728 miles?
Answer by Alan3354(69443)

Can you please explain how to get a answer to this question. Bob and Betty traveled 392 miles in 7 hours.
===
Find the rate.
d = r*t
r = d/t
-----
If they travel at the same rate, how long will it take them to travel 728 miles?
d = r*t
t = d/r